OpenStudy (anonymous):

y=x^2 and y=sinx. Find the area using integration.

OpenStudy (anonymous):

how would you know the interval? or the limit?

OpenStudy (anonymous):

its that old equation again!

OpenStudy (anonymous):

sorry.

OpenStudy (anonymous):

x=0, 0.8767 you need to integrate both functions between these two limits and subtract i'm assuming youe need the area between the two curves

OpenStudy (anonymous):

i'm referring to x^2 = sin x

OpenStudy (anonymous):

yes i do. but you got 0.8767 by using the graph?

OpenStudy (anonymous):

- i got it on wolfram alpha! - the graph yes

OpenStudy (anonymous):

our professor told us to get the value of the x using factorial. how do we do that?

OpenStudy (anonymous):

oh i see - to be honest i dont know chakaron i'll have to check that out factorial or factoring?

OpenStudy (anonymous):

one way would be is to draw the graph to get a rough figure then use trial and error until you get a more accurate answer but that is very tedious. using a calculator would speed it up.

OpenStudy (anonymous):

its factorial. (e.g 5!=1x2x3x4x5) yes. I've tried the graph too but he said (professor) that it wouldn't be accurate.

OpenStudy (anonymous):

For the integration of \[\int\limits_{0}^{1} (4-x2)dx \] what is the answer?

OpenStudy (amistre64):

most trig functions can only be expressed by approximations even when you know the power series you still cant determine its value since there are an infinite number of terms in the polynomial expressed

OpenStudy (anonymous):

so there is no definite answer?

OpenStudy (amistre64):

if you cant determine an exact answer for your bounds, then you will never be able to determine an exact area from it

OpenStudy (amistre64):

your best bet would determine the level of accuracy that you want it to be

OpenStudy (anonymous):

ok. thanks=)

myininaya (myininaya):

its not possible to solve x^2=sinx without using a numerical method such as newton's method and i'm pretty sure there are other methods as well

OpenStudy (amistre64):

numerical methods dont provide an exact solution either :)

myininaya (myininaya):

i mean it is easier to see x^2=sinx when x=0 but the other x you will need numerical approach

myininaya (myininaya):

i know

myininaya (myininaya):

but you can get close to the number

OpenStudy (amistre64):

yes, close is as close gets ;)

OpenStudy (anonymous):

there is a method called linear interpolation - a numerical method which will give you an answer to 4-5 decimal places reasonably quickly

myininaya (myininaya):

so .8767 is pretty close

OpenStudy (anonymous):

i guess so

myininaya (myininaya):

use newton's method and chose intial guess 1 and did 5 iterations and got on the fifth iteration .876741199 the first 4 number on the 4th iteration were the same and i did do one more but i forgot to write it down the first 4 numbers were the same so it seems to be getting pretty close to .8767

OpenStudy (anonymous):

that s good - wolframalpha gave .876727